OpenStudy (anonymous):

Which equation has a single solution of x = 1.2? a)0 = 25x2 + 60x + 36 b)0 = 25x2 – 60x + 36 c)0 = 36x2 + 60x + 25 d)0 = 36x2 – 60x + 25

OpenStudy (anonymous):

@pitamar do you know this

OpenStudy (anonymous):

Remember how we found the roots in the previous problems? we had something of the kind \(f(x) = (x+a)(x+b)\) and from that we knew what values make the function result in 0. Having a single solution means that 'a' and 'b' are equal and therefore we got something of the kind \(f(x) = (x+a)^2\). So try and form each function in the following form: \(f(x) = (x+a)(x+b)\) Then see which ones are capable of having a single solution, and then see if the solution 1.2 causes them to evaluate to 0

OpenStudy (anonymous):

how would i know what is a and b

OpenStudy (anonymous):

you have to transform an expression like 25x^2 + 60x + 36 into something of the form (x+a)(x+b) that is equal to it. You know that kind of stuff?

OpenStudy (anonymous):

ok so how can i get it in that form

OpenStudy (anonymous):

well. if you look at it, it turns that: $$ (x+a)(x+b) = x(x+b) + a(x+b) = x^2 +xb + ax + ab $$If a and b are equal we can go even further: $$ b = a\\ x^2 +xb + ax + ab = x^2 + xa + ax + aa = x^2 + 2ax + a^2 $$ We have to try and reverse the process

OpenStudy (anonymous):

Let's take an example.

OpenStudy (anonymous):

this is a little confusing

OpenStudy (anonymous):

Can you 'open' this? $$(x+1)(x-3) = ?$$

OpenStudy (anonymous):

yea one moment

OpenStudy (anonymous):

how do i reopen it

OpenStudy (anonymous):

$$ (x+1)(x-3) = x(x-3) + 1(x-3) = ... $$

OpenStudy (anonymous):

so i will have to distribute?

OpenStudy (anonymous):

Ye that's what i'm asking you to do right now

OpenStudy (anonymous):

\[x ^{2}\]

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

waiting

OpenStudy (anonymous):

its (x2 +1x)(x2-3)=(x2-3x)+(1x-3)

OpenStudy (anonymous):

well it really is (x2-3x)+(1x-3), but I don't really understand the (x2 +1x)(x2-3) part. Anyway, let's replace the 3 and 1 with 'a' and 'b' and you try again: $$ (x+a)(x+b) = ? $$

OpenStudy (anonymous):

so it would be (x+3) (x+1)

OpenStudy (anonymous):

No, it's a different expression, just like you did with the (x+1)(x-3) I want you to open this expression

OpenStudy (anonymous):

so distribute (x+1)(x-3)?

OpenStudy (anonymous):

That is what you've just distributed. Now distribute this: $$ (x+a)(x+b) = ... $$

OpenStudy (anonymous):

ok (x+a)(x+b)=... x2+ab?

OpenStudy (anonymous):

well not exactly let's see $$ (x+a)(x+b) = x(x+b) + a(x+b) = ... $$

OpenStudy (anonymous):

ok (x2+bx)+ax+ab)?

OpenStudy (anonymous):

ye, exactly. and that's what I wrote above that you thought was confusing

OpenStudy (anonymous):

oh ok lol thanks so whats next now

OpenStudy (anonymous):

a and b can be any numbers. it could be 3... 5.. whatever. it will always end up in that form. Now let's say a = b. Then we have $$ a = b\\ (x+a)(x+b) = (x+a)(x+a) = ...? $$

OpenStudy (anonymous):

so it will be (x2+a)?

OpenStudy (anonymous):

heh. again: $$ (x+a)(x+a) = x(x+a) + a(x+a) = ? $$

OpenStudy (anonymous):

ok (x2+ax)+(ax+a)

OpenStudy (anonymous):

close a*a is what?

OpenStudy (anonymous):

will it just be aa? or just a?

OpenStudy (anonymous):

\(a\cdot a\) (which is the same as writing \(aa\)) is basically \(a^2\)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so its a2

OpenStudy (anonymous):

ye, so (x+a)(x+a) ends up with what?

OpenStudy (anonymous):

(ax)(ax)?

OpenStudy (anonymous):

or (a2)(a2)

OpenStudy (anonymous):

Neither is correct. go over what we did again, then try, ok?

OpenStudy (anonymous):

ok is it (x2)(a2)

OpenStudy (anonymous):

you tend to forget the step $$ (x+a)(x+a) = x(x+a) + a(x+a) $$

OpenStudy (anonymous):

(x2+ax)+(ax+a2)

OpenStudy (anonymous):

Correct. Maybe this visualization will help:  Each side of this square is in length of (x+a). The area of the square (which is literally \((x+a)^2\)) is the sum of the components inside.